Is T: M2,2 → ℝ defined by T(A) =|A| a linear transformation? Provide a proof or counterexample.
Answer #1

i2 (0) Question: Let T:M32. > R be a function feom the set Maza to the Jef R, defined by T(A) = IAL -(1) whese M Mza is the s(2) clearly the given function To Maja >R, defined by T(A) = 1 AL fol every AE Marzo As not a lineal transformation. For exam13) Now by definition of the function I given by (l), we have TIA) o = (1)0) – (0) (0),{by 07] AL T[ 1) = 0 -18) - Tỵ ) Agaj149 have T(4) FT (B) = 0+0 = 0 -o T(A)+. T(B) = 0 (11) Flow equations (10) and (II), we have T(A+B) = 1 I and T(A)+T(B) = 0 .